# Bosonization in $R$-paraparticle Luttinger models

### v2

@@ Abstract
[…]
Copyright attribution to authors.
This work is a submission to SciPost Physics.
License information to appear upon publication.
Publication information to appear upon publication.
Received Date
Accepted Date
Published Date
1 Introduction
@@
Conventionally, particles are categorized as either bosons or fermions; their difference lies primarily with their inherent exchange statistics, where the boson (fermion) wavefunction is symmetric (antisymmetric) under particle exchange. An exception to this dichotomy are anyons [1, 2], existing only in two dimensions [3, 4] and where the particle’s wavefunction pick up an arbitrary phase under an exchange of particles.
@@
Outside anyons, multiple theories of exotic particles that do not follow bosonic or fermionic statistics have also been proposed: all of which have not yet been observed experimentally. Probably the most well-studied are Green’s paraparticles [5], which to this day garners opposing opinions on its actual existence in nature [6, 7, 8]. While superselection rules [9] and no-go theorems [10, 11] tell us that under certain assumptions Green’s paraparticles can always be re-expressed in terms of ordinary bosons and fermions [12, 13], these studies do not explicitly rule out the possibility of them existing in nature. The field has become purely theoretical, with paraparticles being explored through their thermodynamic signatures [14], using first-quantized parastatistics theories [8, 52, 53, 54] and para-manifolds [15], and as potential candidates for dark matter [16, 17, 18]. To date, the only experimental efforts to studying paraparticles are trapped ion simulations of para-bose and para-fermi oscillators [19, 20, 21, 22].
Recently, an alternative formulation was proposed which makes use of generalized commutation relations due to a four-tensor $R$ [23]. In this “$R$-paraparticle” formulation, quasiparticle excitations with parastatistical behavior can be constructed using nonlocal string operators from an underlying set of local spin operators. This approach evades the no-go theorems [10, 11] which assumes that excitations are only created by local operators. Moreover, the superselection rules [9] are inherently satisfied by quasiparticle excitations. This construction from local spin operators opens the possibility of realizing emergent paraparticles in condensed matter systems and in quantum simulations using trapped ions [24]. Furthermore, emergent $R$-paraparticles realized in higher dimensional condensed matter systems enable a novel way of nonlocal secret communication, as proposed in [6], which demonstrates their fundamental physical distinction from ordinary fermions and bosons, and provides an experimental way to observe their exotic exchange statistics. However, the discussions in [23] primarily focus on free paraparticle systems. In contrast, conventional bose and fermi systems exhibit a wealth of collective excitations and emergent phases arising from particle interactions. This naturally raises the question of how interactions might influence the behavior of paraparticles, making it an intriguing direction for further investigation.
The Luttinger model [25, 26, 27, 28], which describes a one-dimensional fermionic gas at low temperatures, plays a significant role in understanding interacting quantum systems. A key feature of this model is bosonization, where fermionic degrees of freedom collectively behave as bosonic modes. This allows the system’s Hamiltonian to be rewritten purely in terms of bosonic operators, making the originally interacting fermionic system exactly solvable [29]. As a result, the model exhibits several non-Fermi liquid properties, including spin–charge separation [30, 31], power-law decay of correlation functions [26, 27, 32], and heightened sensitivity to impurities [33, 34, 35].
@@
In this paper, we study a generalized version of the Luttinger model where we extend the conventional fermionic operators to $R$-paraparticle operators. This approach of extending the operators to $R$-paraparticle operators has recently been explored in the context of para-SYK models [36, 37]. We show that bosonization also occurs for $R$-paraparticles with fermi-like average occupation numbers [14], which we call “$R$-parafermions” in analogy to Green’s paraparticles [5, 9]. This presents a method in which interacting $R$-paraparticle systems can be probed using standard methods of solving bosonic systems [29]. Before proceeding to the remainder of the study, we first clarify that the term “parafermion” should not be confused with generalizations of Majorana fermions [38, 39, 40], which unfortunately share the same terminology in condensed matter literature. For instance, the work in [41] also addresses the bosonization of parafermions; however, the parafermions considered there are those related to Majorana fermions and are not the $R$-parafermions considered in this work.
@@
The paper shall be organized as follows. In Sec. 2, we review the $R$-paraparticle formulation and clarify how $R$-parabosons and $R$-parafermions are classified based on their thermodynamics and exclusion statistics. We show that $R$-parafermions obey Pauli’s exclusion principle or one of its generalizations, prompting us to define $p$-orderedness in the $R$-paraparticle formalism. In Sec. 3, $R$-parafermionic density wave excitations are shown to be bosonic, independent of the number of degrees of freedom of the $R$-paraparticle, and are completely decoupled to flavor waves. However, for systems with two internal degrees of freedom, flavor waves, which are generalizations to spin waves, are bosonic only for specific $R$-parafermion species. More importantly, we show that bosonization only applies at low energies for $R$-paraparticles that satisfy Pauli exclusion principle, as $R$-parafermions obeying higher-order exclusion principles generate partition functions that do not match in the $R$-parafermionic and bosonic bases. These results point to possible experimental realizations of $R$-paraparticle signatures in one-dimensional systems by searching for signs of flavor-charge separation [31] in the absence of magnon excitations [42], as will be discussed in Sec. 4. Finally, we summarize the contents of this paper and provide possible extensions of this work in Sec. 5.
2 R-paraparticles
2.1 Formulation
@@
the second equation being the constant Yang-Baxter equation (YBE) [43, 44, 45]. A paraparticle that is described by a unique $R^{ab}_{cd}$ then satisfies the generalized commutation relations (GCR)
@@
where $\hat{\psi}_{i,a}^{+}$ ($\hat{\psi}_{i,a}$) creates (annihilates) a particle at mode $i$ with an internal or flavor degrees of freedom indexed by $a$. The integer $m$ denotes the maximum number of internal degrees of freedom (which need not be related to any observable [13]) considered in the interaction. The creation and annihilation operators satisfy $\hat{\psi}_{i,a}^{+}=(\hat{\psi}_{i,a})^{\dagger}$ only if $R$ satisfies
@@
The single-mode partition function for an $R$-paraparticle is defined as
@@
where $x\equiv e^{-\beta\epsilon}$, $\beta=1/k_{B}T$, and $\epsilon$ is the $R$-paraparticle energy. $d_{n}$ are nonnegative integers and describe the dimension of the $n$-particle state. Importantly, $d_{n}=0$ implies that the $n$ $R$-paraparticles cannot occupy the same mode at once. The average occupation number then can be calculated as
@@
For the remainder of this paper, we shall reshape the $R$-tensor as an $m$-dimensional matrix $M_{uv}=R^{ab}_{cd}$, where u=m(a1)+bu=m(a-1)+b and v=m(c1)+dv=m(c-1)+d. As an example, $M$ matrices for the $m=2$ case can be written as
@@
2.2 R-parafermions
Analogous to Green’s formulation which generalizes parafermions with integer- and half-integer spins as parabosons and parafermions [5], we also construct a classification scheme for $R$-paraparticles. Suppose the GCR due to $R$ allows at most $n^{\prime}$ particles to occupy a single mode, irrespective of particle flavor. This implies that $d_{n>n^{\prime}}=0$, hence the average occupation number is reduced to
@@ 2.2 R-parafermions
i.e., the occupation number close to absolute zero is proportional to a Heaviside step function $\theta(x)$, and therefore the $R$-paraparticle possesses a fermi surface structure. This motivates us to classify $R$-paraparticles that satisfy Eq. (11b) as $R$-parafermions. Analogously, we show in the Appendix that when $d_{n}\neq 0$ for all $n$, the average occupation of the particle resembles bose statistics, and therefore these $R$-paraparticles are categorized as $R$-parabosons. In Green’s paraparticle formulation, these behaviors were also observed in [14].
The requirement that $d_{n>n^{\prime}}=0$ implies an exclusion principle is at play within the $R$-paraparticle’s GCR. In general, this can be expressed as
@@
where $p=1$ is the usual Pauli exclusion principle. This definition is analogous to Green’s $p$-ordered parafermions, to which there can be at most $p$ particles in the same quantum state [9]. We therefore define a $p$-ordered $R$-parafermions as $R$-paraparticles that satisfy Eq. (12), i.e., allowing at most $p$ particles of the same flavor in a single mode. The maximum number of $R$-paraparticles that can occupy a single mode, $n^{\prime}$, need not be equal to $p$.
For $m=2$, examples of $p=1$ $R$-paraparticles include those that are described by the $M$-matrices
@@
$M_{2}$ $=\begin{pmatrix}-1&0&0&0\\ 0&0&\alpha&0\\ 0&1/\alpha&0&0\\ 0&0&0&-1\end{pmatrix}$, (13b)
@@
where $\alpha\in\mathbb{C}_{\neq 0}$. The $R$-tensor of $M_{1}$ is exactly example 3 of [23] where $n^{\prime}=1$, while both $M_{2}$ and $M_{3}$ have $n^{\prime}=2$. Note that the matrices $M_{2}$ with $|\alpha|\neq 1$ and $M_{3}$ describe non-unitary $R$-tensors, and ordinary fermions are retrieved from $\alpha=-1$ in Eq. (13b). Likewise, $p=2$ cases also exist for $R$-paraparticles with $m=2$, such as those described by
$M_{4}=\begin{pmatrix}0&0&0&\beta\\ 0&-1&0&0\\ 0&0&-1&0\\ 1/\beta&0&0&0\end{pmatrix}$, (13d)
also with $\beta\in\mathbb{C}_{\neq 0}$. This also describes a non-unitary $R$-tensor whenever $|\beta|\neq 1$. The exclusion statistics of $R$-paraparticles described by $M_{4}$ go as $\hat{\psi}_{i,a}^{+}\hat{\psi}_{i,b}^{+}=-\hat{\psi}_{i,a}^{+}\hat{\psi}_{i,b}^{+}=0$ and $\hat{\psi}_{i,a}^{+}\hat{\psi}_{i,a}^{+}\propto\hat{\psi}_{i,b}^{+}\hat{\psi}_{i,b}^{+}\neq 0$ for $\in\{1,2\}, aba\neq b$. This results to $\hat{\psi}_{i,a}^{+}\hat{\psi}_{i,a}^{+}\hat{\psi}_{i,a}^{+}\propto\hat{\psi}_{i,a}^{+}\hat{\psi}_{i,b}^{+}\hat{\psi}_{i,b}^{+}\allowbreak=0$, implying that three particles cannot occupy the same mode at once, and therefore $d_{n>2}=0$.
3 R-parafermion bosonization
Having defined $p$-ordering and the classification of $R$-paraparticles into $R$-parafermions or $R$-parabosons, let us now consider the bosonization of one-dimensional interacting $R$-paraparticles. We define the $R$-paraparticle Luttinger model kinetic Hamiltonian term as

### v3

@@ Abstract
Alternative theories of quantum statistics provide an avenue for exploring novel physics beyond bosons and fermions, yet experimental verification of their existence in nature proves a challenging task. Among these theories, it has recently been suggested that $R$-parastatistics can be realized as quasiparticle excitations in many-body systems. In this paper, we build on this idea by showing that signatures of $R$-parastatistics can be observed as flavor-charge separation in 1D systems. We consider a generalized version of the Luttinger model and show that bosonization persists when the $R$-paraparticles have fermi-surface-like structures. These $R$-parafermions can satisfy generalized exclusion principles beyond conventional Pauli’s. We show that density waves of all $R$-parafermions can always be bosonized, but flavor waves act like bosons only for a certain sublcass of $R$-parafermions. We derive the conditions for bosonization by analyzing the LM spectrum, showing that bosonization applies only to low-temperature systems. Signatures of flavor-charge separation then become apparent as distinct dispersion profiles when we turn on inter-particle interactions. This points to potential observations of flavor-charge separation in 1D systems that host emergent $R$-paraparticles.
preprint: APS/123-QED
I Introduction
@@
Conventionally, particles are categorized as either bosons or fermions; their difference lies primarily with their inherent exchange statistics, where the boson (fermion) wavefunction is symmetric (antisymmetric) under particle exchange. One popular exception to this dichotomy are anyons [1, 2], existing only in two dimensions [3, 4] and where the particle’s wavefunction picks up an arbitrary phase under an exchange of particles.
Figure 1: Summary of main results. (a.) Occupation statistics of different types of $R$-parafermions (solid line) and $R$-parabosons (dashed line). The single-mode partition functions of the latter are obtained using Eq. (44). (b) Partition functions of the noninteracting $R$-paraparticle LMs in the $R$-parafermion and boson basis. (c) Sketch of expected dispersions of $R$-parafermionic quasiparticles obtained from an underlying spinor TG gas.
@@
Outside of anyons, multiple theories of unconventional quantum statistics have also been proposed [5, 6, 7, 8], and although they present novel physical predictions [9, 10, 11, 12], no unambiguous experimental evidence of these unconventional particles has yet been reported. Among these theories is Green’s parastatistics [5], which remains a subject of ongoing discussion concerning its actual occurrence in nature [13, 14]. While superselection rules [15] and no-go theorems [16, 17] tell us that under certain assumptions Green’s paraparticles can always be re-expressed in terms of ordinary bosons and fermions [18, 19], they do not explicitly rule out the possibility of their existence in nature. For a long time, the field has remained purely theoretical, being explored through different approaches such as their thermodynamic signatures [20], under first-quantized theories [14, 21, 22, 23], para-manifold formulations [24], and as potential candidates for dark matter [25, 26]. To date, the only experimental efforts to studying paraparticles are trapped ion simulations of para-bose and para-fermi oscillators [27, 28, 29, 30, 31].
Recently, an alternative paraparticle formulation was proposed that allows the construction of these systems as emergent quasiparticles [32]. This recently introduced $R$-paraparticle framework showed that quasiparticle excitations with parastatistical behavior can be constructed using nonlocal string operators from an underlying set of local spin operators. This approach evades the no-go theorems [16, 17] which assume that excitations are created by local operators. Moreover, the superselection rules [15] are inherently satisfied by quasiparticle excitations. $R$-paraparticles are predicted to have practical applications in quantum communication [33, 34].
This then motivates the realization of emergent $R$-paraparticles, potentially in condensed matter systems or trapped ion simulations [35]. An indirect method of doing this is to look for exotic many-body signatures in the form of collective excitations and emergent phases that are due to the unconventional exchange statistics of the system. It is therefore interesting to ask how interactions would influence the behavior of $R$-paraparticles. For this purpose, we then consider an $R$-paraparticle system described by an interacting Luttinger model (LM) [36, 37, 38, 39]. Conventionally, this describes a 1D fermionic gas at low temperatures. In this work, we are mainly interested in its property known as bosonization, where fermionic degrees of freedom collectively behave as bosonic modes. This allows the system’s Hamiltonian to be rewritten purely in terms of bosonic operators, making the originally interacting fermionic system exactly solvable [40]. As a result, the model exhibits several non-Fermi liquid properties, including spin–charge separation [41, 42].
@@
In this paper, we study a generalized version of the LM where we extend conventional fermionic operators to $R$-paraparticle operators. We show that bosonization also occurs for $R$-paraparticles with a fermi-like average occupation number, which we call $R$-parafermions in analogy to Green’s paraparticles [5, 15, 20]. This presents a method in which interacting $R$-paraparticle systems can be probed using standard methods of solving bosonic systems [40]. However, before we proceed to the remainder of the study, we first clarify that the term “parafermion” should not be confused with generalizations of Majorana fermions [43, 44, 45], which unfortunately share the same terminology in condensed matter literature. More importantly, while the work in [46] also addresses bosonization of parafermions, our work is completely unrelated to the latter.
@@
The paper shall be organized as follows. In Sec. II, we review the $R$-paraparticle formulation and clarify how $R$-parabosons and $R$-parafermions are classified based on their thermodynamics and exclusion statistics, c.f., Fig. 1(a). We show that $R$-parafermions obey Pauli’s exclusion principle or one of its generalizations, which allows us to define $p$-orderedness in the $R$-paraparticle formalism. In Sec. III, we show that all $R$-parafermionic density wave excitations have bosonic properties independent of the number of degrees of freedom of the $R$-paraparticle, and that the density wave operator always commutes with flavor wave operators, which generalizes the spin wave operator. However, for systems with two internal degrees of freedom, flavor waves are bosonic only for specific $R$-parafermion species, hence only shows flavor-charge separation in specific cases. More importantly, we show that bosonization only applies at low energies for $R$-paraparticles, independent of its exclusion principle, as shown in Fig. 1(b). We also show that by introducing interaction, the flavor and density waves of bosonizable $R$-parafermion systems obtain different group velocities. In Sec. IV, we make use of these ideas to lay out a possible experimental design for realizations of $R$-paraparticle signatures in 1D systems by searching for signs of flavor-charge separation [42] in the absence of spinon excitations. For an ideal underlying system which hosts $R$-paraparticle quasiparticles, we discuss in detail spinor Tonks-Girardeau (TG) gases [47] and its spin-charge separation behavior, which we summarize in Fig. 1(c). Briefly, the boson system possesses a dispersion $\epsilon_{k}=-2t\cos(k)$ that is purely due to its charges, since spin-charge separation in TG gases implies $v_{S}=0$ [48]. $R$-parafermionic excitations can arise from the spinor TG gas system, and in the LM limit, results to both nonzero flavor and charge wave velocities. In the presence of interaction between $R$-paraparticles, $\epsilon_{k}$ is altered but $v_{S}$ remains zero, i.e., the spins remain stationary. However for the $R$-parafermions, the flavor and charge velocities become unequal, $v_{C}\neq v_{F}$, which becomes a telltale signature of flavor-charge separation. Finally, in Sec. V, we summarize the contents of this paper and provide possible extensions of this work.
II R-paraparticles
II.1 Formulation
@@
where the latter is the constant Yang-Baxter equation (YBE) [49, 50, 51]. These conditions impose that an $R$-paraparticle wavefunction returns to itself after two consecutive swaps of the same pair of particles. An $R$-paraparticle that is described by $R^{ab}_{cd}$ then satisfies the generalized commutation relations (GCR)
@@
where $\hat{\psi}_{i,a}^{+}$ ($\hat{\psi}_{i,a}$) creates (annihilates) a particle at mode $i$ with an internal or flavor degrees of freedom indexed by $a$. The integer $m$ denotes the maximum number of internal degrees of freedom considered in the interaction. The creation and annihilation operators satisfy $\hat{\psi}_{i,a}^{+}=(\hat{\psi}_{i,a})^{\dagger}$ only if $R$ satisfies
@@
The above unitarity requirement imposes a simplified construction of the orthonormal state space due to $R$. When $R$ is non-unitary (i.e., it does not satisfy Eq. (3)), an alternate description of the Hermitian inner product is required to impose Hermiticity of obervables [32].
@@
The single-mode partition function for an $R$-paraparticle can be defined as
@@
where $x\equiv e^{-\beta\epsilon}$, $\beta=1/k_{B}T$, and $\epsilon$ is the $R$-paraparticle energy. $d_{n}$ are nonnegative integers and describe the dimension of the $n$-particle state space. Importantly, $d_{n}=0$ implies that $n$ $R$-paraparticles cannot occupy the same mode at once. The average occupation number can then be calculated as
@@
In general, different sets of GCRs can result to the same set of $\{d_{n}\}_{n=1}^{\infty}$, hence different $R$-paraparticles can possess similar occupation statistics.
For the remainder of this paper, we shall reshape the $R$-tensor as an $m^{2}\times m^{2}$ matrix $M_{uv}=R^{ab}_{cd}$, where u=m(a1)+bu=m(a-1)+b and v=m(c1)+dv=m(c-1)+d. As an example, $M$ matrices for the $m=2$ case can be written as
@@
II.2 R-parafermions
Consider the case where the exclusion statistics due to $R^{ab}_{cd}$ allows at most $n^{\prime}$ particles to occupy a single mode. This implies that $d_{n>n^{\prime}}=0$, hence the average occupation number is reduced to
@@ II.2 R-parafermions
i.e., the occupation number close to absolute zero is proportional to a Heaviside step function $\theta(x)$, and therefore the $R$-paraparticle possesses a fermi-surface-like structure. This motivates the classification of $R$-paraparticles that satisfy Eq. (11b) as $R$-parafermions. Analogously, we show in the Appendix that when $d_{n}\neq 0$ for all $n$, the average occupation of the particle resembles bose statistics, and therefore these $R$-paraparticles are categorized as $R$-parabosons. A similar classification scheme was also used in Refs. [52, 53], albeit not expounded in terms of the $R$-paraparticles’ occupation statistics. In Fig. 1(a), we present several examples of the occupation statistics of $R$-paraparticles described by different single-mode partition functions. The partition functions $z(x)=(1+x)^{m}$ are representative of $m$-component fermions, while $z(x)=1+mx$ describes $n^{\prime}=1$ $R$-parafermions. $R$-paraboson occupations described by the dashed lines are obtained from $R^{\prime}=-R$, where $R$ describes the $R$-parafermion designated with the same curve color. These occupation behaviors were also predicted in Green’s paraparticle formulation [20].
The requirement that $d_{n>n^{\prime}}=0$ implies an exclusion principle is at play within the $R$-paraparticle’s GCR. In general, this can be interpreted as the operator algebra satisfying
@@
where $p=1$ reduces to Pauli’s exclusion principle. This definition is analogous to Green’s $p$-ordered parafermions, to which there can be at most $p$ particles in the same quantum state [15]. We therefore define a $p$-ordered $R$-parafermions as $R$-paraparticles that satisfy Eq. (12), i.e., allowing at most $p$ particles of the same flavor in a single mode. By definition, we have $n^{\prime}\geq p$.
For the $m=2$ case, examples of $p=1$ $R$-paraparticles include those that are described by the $M$-matrices
@@
$M_{2}$ $=\begin{pmatrix}-1&0&0&0\\ 0&0&\gamma&0\\ 0&1/\gamma&0&0\\ 0&0&0&-1\end{pmatrix}$, (13b)
@@
where $\gamma\in\mathbb{C}_{\neq 0}$. The $R$-tensor defined by $M_{1}$ gives $n^{\prime}=1$ and is exactly Ex. 3 of [32], while both $M_{2}$ and $M_{3}$ have $n^{\prime}=2$. Note that the matrices $M_{2}$ with $|\gamma|\neq 1$ and $M_{3}$ describe non-unitary $R$-tensors, and ordinary fermions are retrieved from $\gamma=-1$ in Eq. (13b). Likewise, $p=2$ cases also exist for $R$-paraparticles with $m=2$, such as those described by
$M_{4}=\begin{pmatrix}0&0&0&\delta\\ 0&-1&0&0\\ 0&0&-1&0\\ 1/\delta&0&0&0\end{pmatrix}$, (13d)
also with $\delta\in\mathbb{C}_{\neq 0}$. This also describes a non-unitary $R$-tensor whenever $|\delta|\neq 1$. The exclusion statistics of $R$-paraparticles described by $M_{4}$ go as $\hat{\psi}_{i,a}^{+}\hat{\psi}_{i,b}^{+}=-\hat{\psi}_{i,a}^{+}\hat{\psi}_{i,b}^{+}=0$ and $\hat{\psi}_{i,a}^{+}\hat{\psi}_{i,a}^{+}\propto\hat{\psi}_{i,b}^{+}\hat{\psi}_{i,b}^{+}\neq 0$ for $\in\{1,2\}, aba\neq b$. This results to $\hat{\psi}_{i,a}^{+}\hat{\psi}_{i,a}^{+}\hat{\psi}_{i,a}^{+}\propto\hat{\psi}_{i,a}^{+}\hat{\psi}_{i,b}^{+}\hat{\psi}_{i,b}^{+}\allowbreak=0$, implying that three particles cannot occupy the same mode at once, and therefore $d_{n>2}=0$.
We end this section with an important note that the definition of order $p$ is not an invariant for $R$-paraparticles. This is because under internal basis transformations, one can map an $R$-paraparticle GCR due to $R^{ab}_{cd}$ to another GCR described by $(R^{\prime})^{ab}_{cd}$, thus following a different set of exclusion statistics. In more detail, we have
$(R^{\prime})^{ab}_{cd}=S^{ab}_{ef}R^{ef}_{gh}(S^{-1})^{gh}_{cd},$ (14a)
or in terms of the $M$-matrix,
$M^{\prime}\tilde{S}=\tilde{S}M,$ (14b)
where
$S^{ab}_{ef}=Q^{a}_{e}Q^{b}_{f}\quad\text{and}\quad\tilde{S}=Q\otimes Q,$ (14c)
with $Q$ an $m\times m$ invertible matrix. Curiously, under this rule, one can show that $M_{2}(\gamma=1)$ is equivalent to $M_{4}(\delta\in\mathbb{C}_{\neq 0})$, i.e., an $R$-paraparticle of order $p=1$ can get mapped to $p=2$. While we have insufficient information on which internal basis an $R$-paraparticle prefers, the important thing to consider in bosonization is the existence of fermi-point-like structures in $R$-parafermions, as we show in the next section.
III R-parafermion bosonization
Having defined $p$-ordering and the classification of $R$-paraparticles into $R$-parafermions or $R$-parabosons, let us now consider the bosonization of 1D interacting $R$-paraparticles. We define the $R$-paraparticle LM kinetic Hamiltonian term as